OpenStudy (studyx3):

Please help, extremely stressed. Will give medal!!!! 4a+5b-6c=2 -3a-2b+7c=-15 -a+4b+2c=-13

OpenStudy (studygurl14):

What's your question?

OpenStudy (studyx3):

I am supposed to solve for a, b, and c

OpenStudy (kevin):

You can find the result using discriminant. Other way, you can use LCM to solve a,b,c value one by one : 4a+5b-6c=2 ----> 12a + 15b - 18c = 6 -3a-2b+7c=-15 --> -12a - 8b + 28c = -60 You'll get new equation contains b and c variables. Do the same for -3a-2b+7c=-15 -a+4b+2c=-13 After you get 2 equations contains just b and c variables, do elimination once again on the first equation and the second equations. After that, you can solve other variables using substitution rule.

OpenStudy (studyx3):

@Kevin I got (-7, -2, -4) for my answer. Is this correct? When plugging them back into one of the original equations (-a+4b+2c=-13), it did not work.

OpenStudy (kevin):

That's not the answer, check again :)

OpenStudy (studyx3):

(9,-2,-4)??

OpenStudy (kevin):

Please, plug them into one of the equations. If it's fit, then it's the answer. If not, then you must check your answer again..

OpenStudy (kevin):

b=-2 is correct c=-4 is correct check again for A

OpenStudy (studyx3):

Okay I finally got the correct answer. Does it matter what equation I use in the beginning? I just don't understand what I was doing wrong..

OpenStudy (kevin):

The equations on the question is used to check your final answer. You must write every new equations carefully. If one of your equations is wrong, then your final answer will be wrong too